# Improving Detection of Watermarked Language Models

### v1

@@ 4 Experimental Results
Tables 1, 2 (Appendix), 9 (Appendix), and 10 (Appendix) show the performance of various watermarking schemes, detectors, and hybrid approaches, when Gemma-7B-instruct is applied to databricks-dolly-15k, human text is taken as the negative class, and the target length is 100 tokens. Unless indicated otherwise, results are discussed under this setting. Many results are deferred to the Appendix.
@@
Among the learnable approaches, we find not much benefit in going from simple LR to a deep MLP or the tree — the gains, if any, are generally small, but there can be losses as well due to potential overfitting on the calibration dataset (more parameters to learn). For example, under Aaronson, MLP improves over LR by an absolute 0.6% (96.8% vs. 97.4%) but Tree loses by 2.7% (94.1%). Broadly, the tree under the hyperparameter setting we explored was more at risk of overfitting and realizing losses than MLP. These insights all suggest that the watermark and non-watermark scores are quite discriminative on their own and do not greatly benefit from excessive learning. Thus, among the learnable options, we recommend simple LR to the practitioner.
@@
Figures 1 and 4 (Appendix) show the effect of entropy on performance. We see that watermarking performance indeed depends on the amount of entropy available. For example, accuracy on the 20% of prompts that afford the least amount of response entropy (whose estimation we described earlier) is about 10% worse in absolute terms than the 20% affording the most entropy. While LLh also improves with entropy, the RoBERTa classifier is both strong and fairly flat, which the hybrid approaches successfully leverage. Concretely, the cascades and logistic regression models with RoBERTa provide a near constant performance in the high 90’s across the entire entropy spectrum. Lastly, hybrid approaches post improvements at all entropy levels.
@@
Tables 17 and 18 (Appendix) show accuracy under the paraphrasing attack when Gemma-7B-instruct is applied to databricks-dolly-15k. We find that paraphrasing effectively removes most of the watermarking signal, as watermark detection is near-random. As a result, the hybrid approaches rely mostly on the non-watermark signal (which alone achieves around 70-80%) and the overall performance boost is minimal. An intriguing exception is RoBERTa, where both LR and MLPs are able to squeeze out additional signal. For example, LR under Aaronson and RoBERTa confers an 8% gain.
@@ A.1 Omitted Discussions
Kuditipudi. We describe the distortion-free algorithm of Kuditipudi et al. [2023] in detail. Consider a secret, finite ordered list of seeds with length . Begin watermarking by first selecting a position in the seed list uniformly at random and then apply the selection rule of Aaronson [2023] with the PRNG seeded to the current value. Advance to the next seed in the list (wrap-around if you are at the end) and repeat. Scoring is performed via a permutation test that evaluates how compatible the query text is with the specific list of seeds used during encoding versus any other random list of seeds of the same length. As the random starting position is not known during scoring, an alignment score based on the Levenshtein distance is formulated that considers alignments of various subsequences of the text and seeds. The proposed method is similar to Aaronson [2023] with the difference of using a fixed list of seeds and a permutation test for scoring. The upside is robustness to attacks; the downside is significantly higher computational cost during scoring. Larger offers more diversity and quality during generation but results in costlier and weaker detection.
@@ Appendix B Omitted Figures
Figures 4 and 5 show, respectively, the effect of entropy and amount of token corruption on accuracy under the Aaronson scheme.
@@ Appendix C Omitted Tables
Tables 2, 3, 4, 5, 6, 7, 8 report accuracy for human negatives and no corruption. Tables 9, 10, 11, 12, 13, 14, 15, 16 show pAUC for human negatives and no corruption. Tables 17 and 18 report accuracy for human negatives under the paraphrasing attack. Lastly, Tables 19, 20, 21, 22 report accuracy for LLM negatives and no corruption.

### v2

@@ 4 Experimental Results
Tables 1, 3 (Appendix), 10 (Appendix), and 11 (Appendix) show the performance of various watermarking schemes, detectors, and hybrid approaches, when Gemma-7B-instruct is applied to databricks-dolly-15k, human text is taken as the negative class, and the target length is 100 tokens. Unless indicated otherwise, results are discussed under this setting. Many results are deferred to the Appendix.
@@
Among the learnable approaches, we find not much benefit in going from simple LR to a deep MLP or the tree — the gains, if any, are generally small, but there can be losses as well due to potential overfitting on the calibration dataset (more parameters to learn). For example, under Aaronson, MLP improves over LR by an absolute 0.6% (96.8% vs. 97.4%) but Tree loses by 2.7% (94.1%). Broadly, the tree under the hyperparameter setting we explored was more at risk of overfitting and realizing losses than MLP. These insights all suggest that the watermark and non-watermark scores are quite discriminative on their own and do not greatly benefit from excessive learning. Thus, among the learnable options, we recommend simple LR to the practitioner. In the Appendix, we analyze the decision boundaries learned by LR for the various non-watermark methods.
@@
Figures 1 and 5 (Appendix) show the effect of entropy on performance. We see that watermarking performance indeed depends on the amount of entropy available. For example, accuracy on the 20% of prompts that afford the least amount of response entropy (whose estimation we described earlier) is about 10% worse in absolute terms than the 20% affording the most entropy. While LLh also improves with entropy, the RoBERTa classifier is both strong and fairly flat, which the hybrid approaches successfully leverage. Concretely, the cascades and logistic regression models with RoBERTa provide a near constant performance in the high 90’s across the entire entropy spectrum. Lastly, hybrid approaches post improvements at all entropy levels.
@@
Tables 18 and 19 (Appendix) show accuracy under the paraphrasing attack when Gemma-7B-instruct is applied to databricks-dolly-15k. We find that paraphrasing effectively removes most of the watermarking signal, as watermark detection is near-random. As a result, the hybrid approaches rely mostly on the non-watermark signal (which alone achieves around 70-80%) and the overall performance boost is minimal. An intriguing exception is RoBERTa, where both LR and MLPs are able to squeeze out additional signal. For example, LR under Aaronson and RoBERTa confers an 8% gain.
@@ A.1 Omitted Discussions
Kuditipudi. We describe the distortion-free algorithm of kuditipudi2023robust in detail. Consider a secret, finite ordered list of seeds with length . Begin watermarking by first selecting a position in the seed list uniformly at random and then apply the selection rule of aaronson with the PRNG seeded to the current value. Advance to the next seed in the list (wrap-around if you are at the end) and repeat. Scoring is performed via a permutation test that evaluates how compatible the query text is with the specific list of seeds used during encoding versus any other random list of seeds of the same length. As the random starting position is not known during scoring, an alignment score based on the Levenshtein distance is formulated that considers alignments of various subsequences of the text and seeds. The proposed method is similar to aaronson with the difference of using a fixed list of seeds and a permutation test for scoring. The upside is robustness to attacks; the downside is significantly higher computational cost during scoring. Larger offers more diversity and quality during generation but results in costlier and weaker detection.
@@ Appendix B Omitted Figures
Figure 4 shows the decision boundary our logistic regression model learns for a specific setting. We see that the learned model consistently puts more weight on the non-watermark detector scores than the watermark ones and this is more pronounced for RoBERTa.
Figures 5 and 6 show, respectively, the effect of entropy and amount of token corruption on accuracy under the Aaronson scheme.
@@ Appendix C Omitted Tables
Table 2 shows the standard deviation of the test accuracy sampling distribution for our hybrid models when the calibration data is bootstrap resampled (100 times) and the test samples are held fixed. Results for each individual detector, 1S, 2S, and LR methods under Aaronson watermarking are depicted when Gemma-7B-instruct is applied to databricks-dolly-15k and human examples are taken as negatives, with a target length of 100 tokens. We see that the standard deviations are low, suggesting low sensitivity to the calibration data and fitting procedure.
| WM | Det | WM Only | Det Only | 1S | 2S | LR |
| LLh | 0.150 | 1.169 | 1.022 | 0.924 | 0.598 | |
| LLR | 0.150 | 0.627 | 1.079 | 0.839 | 0.842 | |
| Aaronson | RoBERTa | 0.150 | 1.233 | 0.649 | 0.647 | 0.051 |
| Binoculars | 0.150 | 0.439 | 0.444 | 0.388 | 0.282 | |
| Radar | 0.150 | 0.000 | 2.456 | 0.343 | 0.268 |
Tables 3, 4, 5, 6, 7, 8, 9 report accuracy for human negatives and no corruption. Tables 10, 11, 12, 13, 14, 15, 16, 17 show pAUC for human negatives and no corruption. Tables 18 and 19 report accuracy for human negatives under the paraphrasing attack. Tables 20, 21, 22, 23 report accuracy for LLM negatives and no corruption.